# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

### v1

@@ A.1. Definitions for Index
Application domain: The specific field or sector in which the system is applied, such as education, healthcare, finance, urban planning, etc.
@@
Social scale: The scale at which the system operates, such as individual, group, organizational, societal, etc.
Theoretical Frameworks: The underlying theories and concepts that support the development and application of the system, which may include theories from cognitive science, social science, computer science, etc.
@@
The metric name
The category to which the metric is assigned |
@@
A.2. User Study Survey
@@
A.3. Data Analysis
| ID | Education Level | UX Research Experience | Coding Experience | AI Literacy Measured by (Ng et al., 2022) | AI Usage in Research |
| P1 | Master’s | 3-4 years | 1-2 years | 4.33 | Research Ideation, Evaluation Plan Development, Coding, Writing, etc. |
| P2 | Master’s | 3-4 years | 5+ years | 5.00 | Research Ideation, Evaluation Plan Development, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P3 | Doctorate | 1-2 years | 3-4 years | 4.67 | Literature Review, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P4 | Master’s | 3-4 years | 5+ years | 4.67 | Research Ideation, Literature Review, Data Analysis, Coding, Writing, etc. |
| P5 | Postdoc | 1-2 years | 3-4 years | 4.00 | Literature Review, Usability Testing, Coding, Writing, etc. |
| P6 | Doctorate | 1-2 years | 3-4 years | 5.00 | Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P7 | Doctorate | 1-2 years | 5+ years | 4.00 | Data Analysis, Coding, Writing, etc. |
| P8 | Doctorate | 3-4 years | 5+ years | 5.00 | Literature Review, Data Analysis, Coding, Writing, etc. |
| P9 | Doctorate | 1-2 years | 3-4 years | 4.00 | Research Ideation, Coding, Writing, etc. |
| P10 | Doctorate | 3-4 years | 3-4 years | 4.00 | Data Analysis, Coding, Writing, etc. |
| P11 | Doctorate | 3-4 years | 5+ years | 3.67 | Research Ideation, Data Analysis, Coding, Writing, etc., Other, please specify |
| P12 | Master’s | 1-2 years | 3-4 years | 5.00 | Research Ideation, Literature Review, Evaluation Plan Development, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P13 | Doctorate | 1-2 years | 3-4 years | 4.00 | Coding, Writing, etc. |
| P14 | Bachelor’s | 3-4 years | 3-4 years | 5.00 | Literature Review, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P15 | Doctorate | 3-4 years | 5+ years | 4.00 | Coding, Writing, etc. |
| P16 | Doctorate | Less than 1 year | 5+ years | 4.33 | Other, debugging front-end UI development |
| P17 | Doctorate | 3-4 years | 5+ years | 4.67 | Other, integrating various LLMs into our embodied robots to support dynamic interactions w/ users |
| P18 | Doctorate | 3-4 years | 5+ years | 5.00 | Research Ideation, Literature Review, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P19 | Master’s | 3-4 years | 5+ years | 4.00 | Prototyping and Design, Coding, Writing, etc. |

### v2

@@ A.1. Definitions for Index
Application domain: The specific field or sector in which the system is applied, such as education, healthcare, finance, urban planning.
@@
Social scale: The scale at which the system operates, such as individual, group, organizational, societal.
Theoretical Frameworks: The underlying theories and concepts that support the development and application of the system, which may include theories from cognitive science, social science, computer science.
A.2. Graph
Table 5 is the table for the graph database nodes and their attributes.
@@ A.2. Graph
The metric name category to which the metric is assigned |
@@
A.3. User Study Survey
@@
A.4. Demographic
Table 6 is the demographic table.
| ID | Education Level | UX Experience | Coding Experience | AI Literacy (Ng et al., 2022) | AI Usage in Research |
| P1 | Master’s | 3-4 years | 1-2 years | 4.33 | RI, EPD, C, W. |
| P2 | Master’s | 3-4 years | 5+ years | 5.00 | RI, EPD, PD, DA, C, W. |
| P3 | Doctorate | 1-2 years | 3-4 years | 4.67 | LR, PD, DA, C, W. |
| P4 | Master’s | 3-4 years | 5+ years | 4.67 | RI, LR, DA, C, W. |
| P5 | Postdoc | 1-2 years | 3-4 years | 4.00 | LR, UT, C, W. |
| P6 | Doctorate | 1-2 years | 3-4 years | 5.00 | PD, DA, C, W. |
| P7 | Doctorate | 1-2 years | 5+ years | 4.00 | DA, C, W. |
| P8 | Doctorate | 3-4 years | 5+ years | 5.00 | LR, DA, C, W. |
| P9 | Doctorate | 1-2 years | 3-4 years | 4.00 | RI, C, W. |
| P10 | Doctorate | 3-4 years | 3-4 years | 4.00 | DA, C, W. |
| P11 | Doctorate | 3-4 years | 5+ years | 3.67 | RI, DA, C, W, Other |
| P12 | Master’s | 1-2 years | 3-4 years | 5.00 | RI, LR, EPD, PD, DA, C, W. |
| P13 | Doctorate | 1-2 years | 3-4 years | 4.00 | C, W. |
| P14 | Bachelor’s | 3-4 years | 3-4 years | 5.00 | LR, PD, DA, C, W. |
| P15 | Doctorate | 3-4 years | 5+ years | 4.00 | C, W. |
| P16 | Doctorate | ¡ 1 year | 5+ years | 4.33 | Other |
| P17 | Doctorate | 3-4 years | 5+ years | 4.67 | Other |
| P18 | Doctorate | 3-4 years | 5+ years | 5.00 | RI, LR, PD, DA, C, W. |
| P19 | Master’s | 3-4 years | 5+ years | 4.00 | PD, C, W. |
A.5. Data Analysis
Table 7 is the data analysis for changes in user perceptions towards the proposed UX plan result.